Solve the differential equation of the following non-exact differential equation. Show complete solution.


(y - xy)dx + xdy = 0


1
Expert's answer
2021-09-30T02:00:38-0400
"xdy=y(x-1)dx"

"\dfrac{dy}{y}=\dfrac{x-1}{x}dx"

Integrate


"\int\dfrac{dy}{y}=\int\dfrac{x-1}{x}dx"

"\ln|y|=x-\ln|x|+\ln C"

"y=\dfrac{Ce^x}{x}"